OpenStudy (anonymous):

What two consecutive numbers add to 257?

OpenStudy (shinalcantara):

Let 'x' be the first number Since they are consecutive numbers then we let 'x+1' be the 2nd number Since their sum would be 257 then you'll have the equation as: x + (x+1) = 257 Solve for x

OpenStudy (anonymous):

So x is 256?

OpenStudy (shinalcantara):

Nope. Solve for x from the derived equation that we have which is x + (x+1) = 257 Removing the parenthesis it will be: x + x + 1 = 257 Transpose 1 to the right x + x = 257 - 1 Combine like terms 2x = 256 Solve for x

OpenStudy (anonymous):

128???

OpenStudy (shinalcantara):

yep. that would be the first number the second number would be x+1

OpenStudy (anonymous):

Is it right? Because 2x128= 256

OpenStudy (anonymous):

What do u mean that the second number is x+1??

OpenStudy (shinalcantara):

since you are looking for 2 numbers which happened to be consecutive and we designate them as 'x' - being the 1st number 'x+1' - the 2nd number then you'll have it as x = 128 x + 1 = 129 Therefore the two consecutive numbers that would add up to 257 would be 128 and 129

OpenStudy (anonymous):

Ahh, ok, i get it now, thanks a lot for the help,

OpenStudy (shinalcantara):

you're welcome

OpenStudy (anonymous):

Thanks again,